# Latent Diffusion Model-Enabled Real-Time Semantic Communication Considering Semantic Ambiguities and Channel Noises

### v1

@@ 7       Determine subsequence $\bm{\tau}=[\tau_{1},\tau_{2},\cdots,\tau_{s}]$;
The distillation training process of EECD is illustrated in Algorithm 3. Additionally, the pretrained latent consistency model can flexibly enhance the perceptual quality of reconstruction by resampling $s-1$ times based on the subsequence $\bm{\tau}=[\tau_{1},\tau_{2},\cdots,\tau_{s}]$ of length $s$, where $\tau_{1}=m$. Consequently, the real-time channel denoising and data reconstruction process based on EECD model is given in Algorithm 4. The advantages of utilizing the enhanced loss term to distill LDM are further elaborated as follows
  • The training of the original LDM cannot optimize the generation of latent space together with the decoder $G_{\bm{\psi}}(\cdot)$. However, the proposed end-to-end consistency loss allows the training objective to no longer be limited to mapping noisy equalization $\bm{y}_{R}$ to the denoised latent space $\hat{\bm{z}}_{0}$, but directly measures the distance between reconstructed data of adjacent data points on the same trajectory.
  • The latent consistency model cannot directly utilize more superior semantic metric functions such as LPIPS, while the EECD based loss effectively eliminates the limitation of only being able to calculate the distance between two latent bottlenecks.
@@ V-A Experimental Setup
Figure 5: Some typical decoded images without/with robust encoder in AWGN 20dB channel and Rayleigh 20dB channel.
@@
1) Dataset: MNIST handwritten digit image dataset is initially considered for evaluating the proposed SemCom system, containing 70,000 grayscale images of handwritten digits from 0 to 9, with 60,000 images for training and 10,000 for testing. Additionally, to validate the performance of out-of-distribution adaptation, the Fashion-MNIST (F-MNIST) dataset is also employed in the evaluation, comprising images of various types of clothing and accessories, with an identical distribution of 60,000 training and 10,000 testing images. The resolution for both MNIST and F-MNIST datasets are uniformly resized to 32×32. Furthermore, the animal face high quality (AFHQ) dataset [65] is also selected to verify the effectiveness, including a total of 15,000 RBG images across three categories: dogs, cats, and wild animals, where 4,500 images of dogs are used for training and the remaining 500 dog images, along with 500 cat images, are used to test the proposed method, with the resolution resized to 192×192. Lastly, the DIV2K high-quality RGB image dataset [66] is also considered for SemCom tasks, encompassing 800 diverse training images, 100 validation images, and 100 test images, with the resolution resized to 256×256.
@@
2) Baseline Method: Three distinct implementations of communication systems are utilized for comparison to demonstrate the superiority of the proposed SemCom system in terms of semantic reconstruction performance and real-time processing capabilities. The first is a combination of the state-of-the-art traditional image compression method JPEG2000 [10] with the error correction technique low-density parity check (LDPC) [67], denoted as JPEG2000+LDPC, which is a common method for image transmission in practical image communications. The second is the widely recognized VAE-based JSCC method [6], where joint source-channel training effectively mitigates the adverse effects of unreliable channels. The final approach is the multi-step VE-based LDM (VE-LDM) [40], which also serves as the teacher model for subsequent knowledge distillation.
3) Performance Metrics: The metrics for model evaluation can broadly be categorized into two types. The first category encompasses traditional bit/symbol accuracy metrics, including measures such as mean squared error (MSE) and peak SNR (PSNR). The second category consists of semantic or human-perceptual metrics that warrant increased attention within the context of SemCom, which for the transmission of images includes the multi-scale structural similarity index measure (MS-SSIM) [15] and the pretrained VGG-based LPIPS [64].
@@
4) CSI Condition: For the wireless channels, three distinct channels are taken into consideration, including the AWGN channel ($K=\infty$), the Rayleigh channel ($K=0$), and the Rician channel ($K=1$). Regarding the noise level of the channels, noise with SNRs ranging from 0dB to 20dB is contemplated for testing the performance of various methods under different SNR conditions. Lastly, for the channel bandwidth ratio (CBR), CBRs from 0.01 to 0.05 are implemented on DL models trained by AFHQ and DIV2K datasets, while for the MNIST dataset, only CBRs of 1/16 and 1/3 have been realized.
@@
5) Simulation Environment and Hyperparameters: The simulations are conducted using Python 3.8.19 and CUDA-accelerated PyTorch 2.3.0 on a computer equipped with an i5-13600KF CPU operating at 3.50GHz, 32 GB of RAM, and an NVIDIA GeForce RTX 4070 GPU. Unlike the data generation task, the training for the wireless channel denoising task can follow a shorter time schedule. Consequently, the total length of the forward process for the LDM is set to $N=100$, with the variance starting point at $t_{1}=\varepsilon=0.002$ and the endpoint at $t_{N}=T=2$, and $\rho=7.0$. It is noteworthy that the image data has been transformed into values ranging from -1 to 1 with both variance and mean of 0.5, while the latent vector comprises data with a mean of 0 and a variance of 1. Furthermore, the learning rate during training is established at 1e-4, with an initial decay rate of 0.95 and a decay rate of 0.99993 for the student model.
@@ V-B Robustness to Data Inaccuracies
As stated in Section III-C, The encoder parameters can be updated via augmented learning based on the obtained semantic errors $\bm{\delta}$, ultimately endowing the DL-based SemCom system with robustness against various types of outliers. For the MNIST, AFHQ, and DIV2K datasets, encoders pretrained are updated with a learning process at an error level of $\left\|\bm{\delta}\right\|_{p}/H=0.3$, where $H$ denotes the image resolution. Following the update, several prototypical image datasets are employed to test the robust encoder’s efficacy in effectively countering data inaccuracies. Fig. 5 illustrates the impact of semantic errors with levels of 0.5 and 0.4 superimposed on the original data under AWGN 20dB and Rayleigh 20dB channel conditions, respectively. It is readily observed that the source data with added semantic errors still bear minimal semantic differences from the original data to the human visual perception. However, when the original SemCom system, without a robust encoder, transmits this contaminated data, the decoded output can result in significant semantic errors, such as the images of dog faces under AWGN 20dB channel conditions and handwritten digits and dog faces under Rayleigh 20dB channel conditions, and might also lead to extensive artifacts in reconstructed images as seen in the DIV2K dataset. Fortunately, the introduction of the robust encoder successfully overcomes semantic ambiguities that may arise from semantic errors or other types of outliers, ensuring that the decoded data at the receiver still carries the correct semantic information.
@@
MNIST PSNR (dB) 16.54/18.50 11.88/18.52 8.33/18.60 5.96/16.36 5.16/12.42 6.17/8.16 7.52/11.35 10.05/14.91 12.30/17.10
SSIM (dB) 10.65/13.32 5.72/13.56 2.77/13.19 1.18/10.56 0.76/6.79 1.40/3.52 2.21/5.15 4.01/8.60 5.77/11.18
MSE 0.022/0.014 0.065/0.014 0.147/0.013 0.253/0.023 0.304/0.057 0.241/0.153 0.177/0.073 0.099/0.032 0.059/0.019
AFHQ PSNR (dB) 22.31/22.51 18.94/22.14 15.34/21.58 12.82/21.38 9.72/20.58 13.56/15.24 18.45/19.19 21.20/21.89 22.32/21.19
MS-SSIM (dB) 19.82/20.68 13.75/20.65 9.34/20.28 6.56/19.11 3.81/17.76 6.76/9.86 13.42/14.27 17.92/18.44 20.27/20.40
LPIPS 0.160/0.152 0.211/0.157 0.302/0.158 0.410/0.172 0.531/0.180 0.475/0.348 0.246/0.226 0.175/0.172 0.154/0.151
DIV2K PSNR (dB) 23.60/23.71 18.62/23.17 14.55/22.70 10.37/22.01 8.99/21.65 11.87/14.27 16.54/18.66 20.87/21.24 22.20/22.41
MS-SSIM (dB) 16.19/16.49 10.20/16.28 8.22/16.01 6.01/15.88 4.94/15.19 5.66/8.39 11.09/12.73 13.68/15.11 15.35/16.22
LPIPS 0.122/0.121 0.208/0.129 0.325/0.133 0.442/0.147 0.560/0.157 0.512/0.297 0.261/0.237 0.184/0.160 0.131/0.128
@@
To maintain generality, the results of multiple performance tests for various outlier types and levels across different datasets are documented in Table I. Specifically, the CBR is fixed at approximately 0.02 and the test CSI conditions vary in accordance with Section V-A. It is not difficult to observe that the robust encoder, despite only being augmentedly trained at a semantic error level of 0.3, also maintains robustness compared to the original encoder under other semantic error levels and low-SNR noise contamination, thereby enhancing the quality of decoded data when source data is subjected to semantic errors or noises. Furthermore, evaluation metrics such as PSNR/SSIM/MS-SSIM can be improved by several times or even an order of magnitude, while MSE/LPIPS can be reduced by several times or even by an order of magnitude.
V-C Out-of-Distribution Adaptation
@@
As described in Section III-D, the proposed SemCom system employs a lightweight, single-layer adapter at the transmitter for rapid one-shot learning and transforms the latent space at the receiver, thereby enabling the DL-based SemCom system to adapt to out-of-distribution data or enhance decoding quality. Specifically, subsets of clothing images from the F-MNIST dataset and cat images from the AFHQ dataset were utilized to validate the efficacy of the adapter. As illustrated in Fig. 6 and Fig. 7, without the adapter enabled, a SemCom system pretrained with a particular type of data would decode data at the receiver that more closely resembles that specific type of semantic information, leading to severe semantic ambiguity. However, the adapter situated before the generator can swiftly overcome this bottleneck, producing data that is essentially consistent with the original semantics of the transmitted data. Additionally, the original training DL model underperformed on certain test data from the DIV2K dataset, with decoded data exhibiting partial errors. The adapter also enhances communication quality in such instances, eliminating artifacts in the images.
@@
Figure 6: Some typical decoded images without/with adaptor in AWGN 20dB channel.
Figure 7: Some typical decoded images without/with adaptor in Rayleigh 20dB channel.
@@
Figure 8: The performance metrics’ curves when do one-shot learning.
@@
The evolution of performance metrics during the one-shot learning process for the three datasets is depicted in Fig. 8. Evidently, after approximately only 20 epochs, the metrics of decoded data with adapters can be swiftly ameliorated to ideal values, thereby diminishing semantic ambiguities. To ensure generality, numerical experiments are also conducted to corroborate the effectiveness of the proposed adaptive strategy in enhancing SemCom performance and mitigating semantic ambiguity, with the results presented in Table II. Notably, as evidenced by the semantic evaluation metrics SSIM/MS-SSIM and LPIPS, the incorporation of adapters substantially augments the receiver’s out-of-distribution adaptation and reconstruction capabilities under certain constraints on image categories, preventing the emergence of semantic ambiguities.
@@
Dataset Metrics PSNR (dB) SSIM/MS-SSIM (dB) MSE/LPIPS
@@
Figure 9: Some typical decoded images with received latent space signals denoised by different approaches in AWGN 10dB channel and Rayleigh 10dB channel.
Figure 10: Semantic metrics (MS-SSIM and LPIPS) of JPEG2000+LDPC, Deep JSCC, VE-LDM, and EECD methods under different SNRs and channel states within AFHQ dataset.

### v4

@@ 4       Compute $\bm{z}\leftarrow E_{\bm{\phi}}(\bm{x})$ and transmitted $\bm{z}_{R}$;
where $\tilde{\bm{z}}^{\bm{\theta}}_{t_{n}}$ denotes the predicted data point at $t=t_{n}$. This is also called denoising diffusion implicit model (DDIM) [22] and every step’s NFE equals to 1. However, the actual values of the difference $\frac{\bm{z}_{t_{n+1}}-\bm{z}_{t_{n}}}{t_{n+1}-t_{n}}$ are closer to the derivatives between $\bm{z}_{t_{n+1}}$ and $\bm{z}_{t_{n}}$, rather than the dervative at $\bm{z}_{t_{n+1}}$. To this end, the Heun solver in EDM is adopted [58], which is denoted by
where $\bm{z}_{t_{n}}$ on the right side of Eq. (27) can be approximated by $\tilde{\bm{z}}^{\bm{\theta}}_{t_{n}}$ in Eq. (26). Consequently, the estimation of $\bm{z}_{t_{n}}$ is given by
where the NFE equals to 2. According to the definition of the consistency function, the function $\bm{f}_{\hat{\bm{\theta}}}(\bm{z}_{t},t)$ should have the same output for adjacent data points $(\bm{z}_{t_{n+1}},t_{n+1})$ and $(\hat{\bm{z}}^{\bm{\theta}}_{t_{n}},t_{n})$ on the same reverse trajectory, i.e., the loss of the consistency model is
where $\hat{\bm{\theta}}^{-}$ denotes the running average of the past values of $\hat{\bm{\theta}}$ during optimization.
Nonetheless, the goal of wireless SemCom is to accurately reconstruct the transmitted data in real-time manner in the receiver side. Inspired by that, the consistency distillation (CD) loss in Eq. (29) can be changed to the loss of EECD, which is given by
where $\bm{d}(\cdot,\cdot)$ is denoted by Euclidean distance for non-images datasets, and structural similarity index measure (SSIM) or learned perceptual image path similarity (LPIPS) [60] based distance for image datasets.
The distillation training process of EECD is illustrated in Algorithm 3. As depicted in Fig. 4, the student model $\bm{f}_{\hat{\bm{\theta}}}(\cdot,\cdot)$ updates its parameters $\hat{\bm{\theta}}$ and $\hat{\bm{\theta}}^{-}$ through gradient descent by minimizing the perceptual loss $\mathcal{L}_{EECD}$ between the decoded data $(\bm{x}_{t_{n+1}},t_{n+1})$ and $(\hat{\bm{x}}^{\bm{\theta}}_{t_{n}},t_{n})$ corresponding to the original latent space data point $(\bm{z}_{t_{n+1}},t_{n+1})$ and the next point $(\hat{\bm{z}}^{\bm{\theta}}_{t_{n}},t_{n})$ in the reverse diffusion process predicted by the diffusion (teacher) model. This process ensures the consistency between denoised signals by direct mapping to $\bm{z}_{\varepsilon}$ along the same ODE trajectory. Additionally, in the sampling phase, the pretrained latent consistency model can flexibly enhance the perceptual quality of reconstruction by resampling $s-1$ times based on the subsequence $\bm{\tau}=[\tau_{1},\tau_{2},\cdots,\tau_{s}]$ of length $s$, where $\tau_{1}=m$. Consequently, the real-time channel denoising and data reconstruction process based on EECD model is given in Algorithm 4. The advantages and contributions of the proposed LDM approach are further elaborated as follows:
  • VE, SDE, and PF-ODE are utilized to model the LDM and wireless channel denoising processes. The advantage of this novel approach lies in its clearer interpretation of physical channels, making it more intuitive and capable of accommodating various channel conditions.
  • The training of the original LDM cannot optimize the generation of latent space together with the decoder $G_{\bm{\psi}}(\cdot)$. However, the proposed end-to-end consistency loss allows the training objective to no longer be limited to mapping received noisy signal $\bm{y}_{R}$ after equalization to the denoised latent space $\hat{\bm{z}}_{0}$, but directly measures the distance of adjacent data points on the same trajectory.
  • The EECD based loss effectively eliminates the limitation of only being able to calculate the Euclidean distance between two latent bottlenecks and multi-step diffusion processes. Consequently, the latent consistency model can directly utilize more superior semantic metrics such as LPIPS to enhance perceptual quality.
@@ V-A Experimental Setup
Figure 5: Some typical decoded images without/with robust encoder under AWGN and Rayleigh channel. The SNR is 20dB.
@@
1) Dataset: MNIST handwritten digit image dataset is initially considered for evaluating the proposed SemCom system, containing 60,000 images for training and 10,000 for testing. Additionally, to validate the performance of out-of-domain adaptation, the Fashion-MNIST (F-MNIST) dataset is also employed in the evaluation, comprising images of various types of clothing and accessories, with an identical distribution of 60,000 training and 10,000 testing images. The resolution for both MNIST and F-MNIST datasets are uniformly resized to 32×32. Furthermore, the animal face high quality (AFHQ) dataset [61] is also selected to verify the effectiveness of the proposed method, including a total of 15,000 RBG images across three categories: dogs, cats, and wild animals, where 4,500 images of dogs are used for training and the remaining 500 dog images, along with 500 cat images, are used to test the proposed method, with the resolution resized to 192×192. Lastly, the DIV2K high-quality RGB image dataset [62] is also considered for SemCom tasks, encompassing 800 diverse training images, 100 validation images, and 100 test images, with the resolution resized to 256×256.
@@
2) Baseline Method: Four distinct implementations of communication systems are utilized for comparison to demonstrate the superiority of the proposed SemCom system. The first is a combination of the state-of-the-art traditional image compression method JPEG2000 [8] with the error correction technique LDPC [43], denoted as JPEG2000+LDPC. The second is the widely recognized CNN-based deep JSCC method [6], where joint source-channel training effectively mitigates the adverse effects of unreliable channels. The third is the multi-step VE-based LDM (denoted as VE-LDM), and the fourth is accelerated DDIM, which involves 2-step sampling [38], both belonging to the DM-aided approaches.
[…]
@@
4) CSI Condition: For the wireless channels, three distinct channels are taken into consideration, including the AWGN channel ($K=\infty$), the Rayleigh channel ($K=0$), and the Rician channel ($K=1$). Regarding the noise level of the channels, noise with SNRs ranging from 0 dB to 20 dB is contemplated for testing the performance of various methods under different SNR conditions. The channel bandwidth ratio (CBR) is defined as $\textrm{CBR}=k/(H\times W\times C)$, where $H$, $W$, and $C$ are the height, width (resolution) and colour channel of images, and usually $H$=$W$. CBR is also an exceedingly crucial metric in SemCom, defining the demand for communication resources [13, 36]. For this reason, CBRs from 0.01 to 0.05 are implemented on DL models trained by AFHQ and DIV2K datasets, while for the MNIST dataset, due to its low resolution, only the DL with 1/16 CBR and the JPEG2000+LDPC with 1/3 CBR have been realized.
@@
5) Simulation Environment and Hyperparameters: The simulations are conducted using Python 3.8.19 and CUDA-accelerated PyTorch 2.3.0 on a computer equipped with an i5-13600KF CPU operating at 3.50GHz, 32 GB of RAM, and an NVIDIA GeForce RTX 4070 GPU. In the encoder and decoder parts, $G_{\bm{\psi}(\cdot)}$ and $E_{\bm{\phi}^{\prime}}(\cdot)$ each contain 7 transposed convolution layers and 6 convolution layers. The training for the wireless channel denoising task can follow a shorter time schedule. Consequently, the total length of the forward process for the LDM is set to $N=100$, with the variance starting point at $t_{1}=\varepsilon=0.002$ and the endpoint at $t_{N}=T=2$, and $\rho=7.0$. Furthermore, the learning rate during training is established at 1e-4, with an initial decay rate of 0.95 and a decay rate of 0.99993 for the student model.
6) Training, Deployment and Testing: During the training and deployment phases, first, the convolutional WGAN and VAE are trained sequentially or the Algorithm 5 is jointly trained and deployed over rate-limited channels; then, the parameters of the trained convolutional VAE will be fine-tuned into a robust encoder in a self-supervised learning manner following the steps of Algorithm 1; finally, the learning of the parameters of the diffusion model is conducted end-to-end according to the denoising EECD strategy in Algorithm 3. In the testing phase, LDM denoises the received equalized signals according to Algorithm 4, and when the reconstruction error exceeds a threshold, the Algorithm 2 will be activated to adjust latent vector $\bm{z}$ for the low-precision data.
@@ V-B Robustness to Data Inaccuracies
As stated in Subsection III-B, the encoder parameters can be updated via augmented learning based on the obtained semantic errors $\bm{\delta}$. For the MNIST, AFHQ, and DIV2K datasets, pretrained encoders are updated with a learning process at an error level of $\left\|\bm{\delta}\right\|_{p}/H=0.3$. Following the update, several prototypical image datasets are employed to test the robust encoder’s efficacy in effectively countering data inaccuracies. Fig. 5 illustrates the impact of semantic errors with levels of 0.5 and 0.4 superimposed on the original data under AWGN channel and Rayleigh channel at SNR of 20 dB, respectively. It is readily observed that with the proposed robust encoder, the source data with added semantic errors still bear minimal semantic differences from the original data to the human visual perception. However, when the original SemCom system, without a robust encoder, transmits this contaminated data, the decoded output can result in significant semantic errors, as shown in Fig. 5 for the example images in MNIST and AFHQ datasets, and might also lead to extensive artifacts in reconstructed images as seen for the DIV2K dataset. Fortunately, the introduction of the robust encoder successfully overcomes semantic ambiguities that may arise from the contaminated data due to cyber attacks or other types of outliers, ensuring that the decoded data at the receiver still carries the correct semantic information.
@@
MNIST PSNR (dB)$\uparrow$ 16.54/18.50 11.88/18.52 8.33/18.60 5.96/16.36 5.16/12.42 6.17/8.16 7.52/11.35 10.05/14.91 12.30/17.10
SSIM (dB)$\uparrow$ 10.65/13.32 5.72/13.56 2.77/13.19 1.18/10.56 0.76/6.79 1.40/3.52 2.21/5.15 4.01/8.60 5.77/11.18
[…]
AFHQ PSNR (dB)$\uparrow$ 22.31/22.51 18.94/22.14 15.34/21.58 12.82/21.38 9.72/20.58 13.56/15.24 18.45/19.19 21.20/21.89 22.32/21.19
MS-SSIM (dB)$\uparrow$ 19.82/20.68 13.75/20.65 9.34/20.28 6.56/19.11 3.81/17.76 6.76/9.86 13.42/14.27 17.92/18.44 20.27/20.40
[…]
DIV2K PSNR (dB)$\uparrow$ 23.60/23.71 18.62/23.17 14.55/22.70 10.37/22.01 8.99/21.65 11.87/14.27 16.54/18.66 20.87/21.24 22.20/22.41
MS-SSIM (dB)$\uparrow$ 16.19/16.49 10.20/16.28 8.22/16.01 6.01/15.88 4.94/15.19 5.66/8.39 11.09/12.73 13.68/15.11 15.35/16.22
[…]
@@
To maintain generality, the results of multiple performance tests for various outlier types and levels across different datasets are documented in Table I. Specifically, the CBR is fixed at approximately 0.02 and the test CSI conditions vary in accordance with Subsection V-A. It is not difficult to observe that the robust encoder, despite only being updated at a semantic error level of 0.3, still maintains robustness compared to the original encoder under other semantic error levels and low-SNR noise contamination, thereby enhancing the quality of decoded data when source data is subject to semantic errors or noises. Furthermore, evaluation metrics such as PSNR/SSIM/MS-SSIM can be improved by several times or even an order of magnitude, while MSE/LPIPS can be reduced by several times or even by an order of magnitude.
V-C Out-of-Domain Adaptation
@@
As described in Subsection III-C, the proposed SemCom system employs a lightweight, single-layer adapter at the transmitter for rapid one-shot learning and transforms the latent space at the receiver, thereby enabling the DL-based SemCom system to adapt to out-of-domain data or enhance decoding quality. Specifically, subsets of clothing images from the F-MNIST dataset and cat images from the AFHQ dataset were utilized to validate the efficacy of the adapter. As illustrated in Fig. 6, without the adapter enabled, a SemCom system pretrained with a particular type of data would decode data at the receiver that more closely resembles that specific type of semantic information, leading to severe semantic ambiguity. However, the adapter situated before the generator can swiftly overcome this shortcoming, producing data that is essentially consistent with the original semantics of the transmitted data. Additionally, the original training DL model underperformed on certain test data from the DIV2K dataset, with decoded data exhibiting partial errors. The adapter also enhances communication quality in such instances, eliminating artifacts in the images.
@@
Figure 6: Some typical decoded images without/with adaptor in AWGN and Rayleigh channel. The SNR is 20 dB.
@@
Figure 7: The performance metrics’ curves when doing one-shot learning to update the parameters of $\bm{g}_{\bm{\omega}}(\cdot)$.
@@
The evolution of performance metrics during the one-shot learning process for the three datasets is depicted in Fig. 7. Evidently, after approximately only 20 epochs, the metrics of decoded data with adapters can be swiftly ameliorated to ideal values, thereby diminishing semantic ambiguities. To ensure generality, numerical experiments are also conducted to corroborate the effectiveness of the proposed adaptive strategy in enhancing SemCom performance and mitigating semantic ambiguity, with the results presented in Table II. Notably, as evidenced by the semantic evaluation metrics SSIM/MS-SSIM and LPIPS, the incorporation of adapters substantially augments the receiver’s out-of-domain adaptation and reconstruction capabilities under certain constraints on image categories, preventing the emergence of semantic ambiguities.
@@
[…]
@@
Figure 8: Some typical decoded images with received signals denoised by different approaches under AWGN and Rayleigh channel. The SNR is 10 dB.
Figure 9: Semantic performance metrics of JPEG2000+LDPC, Deep JSCC, DDIM, VE-LDM, and EECD methods under different SNRs and channel states within AFHQ dataset.